what is the answer step bye step
x-3y=-3
5x-3y=9

Answer by rfer(12678)

x=3y-3
5(3y-3)-3y=9
15y-15-3y=9
12y=24
y=2
--------------------
x-3(2)=-3
x-6=-3
x=4
------------------
5(3)-3(2)=9
15-6=9